A loan of $1465 taken out on June 7 requires three payments. The first payment is due on July 7 The second payment is twice as large as the first payment and is due on August 20. The final payment, due on November 4, is three times as large as the first payment. If the focal date is June 7, what is the size of each of the three payments at an interest rate of 4.5%? The first payment is $255.15, the second payment is $ 510.31, and the third payment is $ 765.46 Round the final answers to the nearest cent as needed. Round all intermediate values to six decimal places as needed.)
media%2Fede%2Fedee6f58-38a4-4b66-89fd-fa
media%2F862%2F86263d9a-853b-404d-a0a4-69
Answer #1
  1. Here P= $5767, r= 28/3%= 9.33%, t= 13/3 years

Hence, I = Prt = 5767 x 28/3 % x 13/3 years

= 2099188/900 = 2332.431111

Hence Interest earned = $2,332.431111

2. Here I= $55.28, t= 162 Days = 162/365 years and r = 4.72% p.a.

Hence, using I=Prt formula,

We find that, 55.28 = P x 4.72% x 162/365

Or 55.28 = P x 764.64/ 36500

Hence P = 36500 x 55.28/764.64 = 2638.784264

Hence Deposit Amount = $2638.784264